# Instructions for Authors: Title in Title Case

### v1

@@ A.2 Proof of Theorem 1
Appendix B Appendix: Experiment Details
B.1 Details of Dataset Settings
@@
Synthetic Dataset: We generated the candidate context $\bm{s}_{I}$ with context dimension $d=10$ as a single context (behavior) for any step of an episode. At step $b$ of $n$-th episode in OL-data phase, candidate context set $\mathcal{S}_{n,b}\subseteq{\mathbb{R}^{d}}$ was drawn from a Gaussian distribution $\mathcal{N}(\mu_{s}\bm{1}_{d},\sigma_{s}^{2}\bm{I}_{d})$, where the means of candidate contexts were $\mu_{s}\in\left[1:-0.4:-2.6\right]$, and the standard deviation was $\sigma_{s}^{2}={0.05}^{2}$. To simulate out-of-distribution(OOD) scenario, the mean of the first candidate context in the BT-data phase was set to $1.4$. We set the observed reward of a given candidate context $\bm{s}_{I_{n,b}}\in\mathcal{S}_{n,b}$ at step $b$ in $n$-th episode as a sigmoid function $\mathrm{sigmoid}(\left\langle\bm{w}_{\mathrm{reward}},\bm{s}_{I_{n,b}}\right\rangle)$, where each element of $\bm{w}_{\mathrm{reward}}\in\mathbb{R}^{d}$ was sampled from a Gaussian distribution as $\mathcal{N}(0.1,{0.01}^{2})$. To test the robustness of the model and simulate random selections of actions made by users, we also added a zero-mean Gaussian noise $\mathrm{std}\sim\mathcal{N}(0,\epsilon^{2}),\epsilon\in\{0.03,0.07,0.1\}$ in the reward, so final reward was set to be $\mathrm{sigmoid}(\left\langle\bm{w}_{\mathrm{reward}},\bm{s}_{I_{n,b}}\right\rangle)+\mathrm{std}$. We also made adjustments for the ablation study period. Contradictory data was introduced as duplicated data: we cloned the first k episodes’ candidate context sets for the second k episode’s candidate context sets, and design k with the name of duplicated quantity $\mathrm{dup}\in\left[0:1:9\right]$. Total number of conflict executed actions (e.g. expert select action 1 at step $b$ of $n$-th episode, but select action 2 at step $b$ of $(n+k)$-th episode.) of the contradictory data was ascending when we increased $\mathrm{dup}$. We also controlled the training rate for all baselines and IBCB: we set the end episode of the exposed expert’s behavioral evolution history logs with constraint end rate $\mathrm{ce\_rate}\in\left[0,1\right]$ ($e.g.$ if $\mathrm{ce\_rate}=0.5$. we use the first $\left\lfloor 0.5N\right\rfloor$ episodes for baselines and IBCB training).
@@
Real-World ML-100K Dataset: We used MovieLens 100K (ML-100K) dataset as our real-world data experiment dataset, which collected 5-star movie ratings from a movie recommendation service, including 943 unique users and 1682 unique movies(items). ML-100K dataset also provided various side information of both users and items. We split ML-100K dataset into OL-Data and BT-data according to the timestamps, where first half was treated as OL-data and the last half was BT-data. For reward setting, we treated the 4-5 star ratings as positive feedback(labeled with 1), and 1-3 star rating as negative feedback (labeled with 0). Through feature embedding [Wang et al., 2019] and principal component analysis (PCA), dimension of each [user, item] pair (candidate context, i.e., behavior) was reduced to 50, $i.e.,\ d=50$. For the candidate context set, we retained the original item for each interaction and then randomly sampled $9$ extra items from the entire item set (exclude the original item), so the number of candidates is $M=10$. In real recommendation, user-item interaction is sparse, meaning that not all of the items can be shown to the user and have feedback [Zhang et al., 2021, 2022]. To overcome this issue, we trained two Matrix Factorization (MF) [Koren et al., 2009] models using OL-data (denotes MF-OL) and BT-data (denotes MF-BT) respectively for different baselines and IBCB model to serve as the simulated real environment. AUCs of the two MF models were both over $90\%$, which assures that the simulated environment can provide nearly realistic feedbacks of users. At step $b$ of $n$-th episode of both OL-data and BT-data phase, the simulated environment (MF-OL/MF-BT) received a selected candidate context $\bm{s}_{I_{n,b}}$ and a context set $\mathcal{S}_{n,b}$ consisted of all candidate contexts in this step, getting all candidate context’s reward and sort out the candidate context $\bm{s}_{I_{\mathrm{max}}}$ with max reward $r_{I_{\mathrm{max}}}$. If selected candidate context $\bm{s}_{I_{n,b}}$ do not have the biggest reward compared to $\bm{s}_{I_{\mathrm{max}}}$, $i.e.,r_{I_{n,b}}<r_{I_{\mathrm{max}}}$, reward feedback would be directly set as 0, otherwise would be set as $r_{I_{\mathrm{max}}}$. If $r_{I_{\mathrm{max}}}>0.5$, then the final reward feedback would be 1, otherwise would be 0. Adjustments for the ablation study period was also made similar with the synthetic dataset. Since duplicated data will not appeared in a large scale in real world, we just controlled the training rate similarly as that in synthetic dataset’s setting.
@@
B.2 Training Details
We have run all of the experiments on the server equipped with 16 Intel Xeon E5-2630 v4@2.20GHz cores, and during training, all of the baselines and IBCB only uses the CPU. Note that all baselines and IBCB were training on expert behavioral evolution history of Online Learning. Detailed settings of baselines and IBCB are shown as follows.
B.2.1 Original Expert Policy Settings

### v2

@@ 4.1.1 Proof of Theorem 1
4.2 IBCB’s Formulation
We will illustrate the formulation using two actions as an example, which can be easily extended to accommodate multiple actions by adding additional constraints. Specifically, given $\mathcal{S}_{n+1,b}=\{\bm{s}_{1},\bm{s}_{2}\}$, =$\mathcal{I}=\{1,2\}$, if action 1 is executed at step $b$ in the $(n+1)$-th episode, we have
which is equivalent to
where $\bm{s}_{1},\bm{s}_{2}\in\mathcal{S}_{2,b},\forall b\in[B].$ More generally, for all $n\in[N-1]$, given a candidate action set $\mathcal{S}_{n+1,b}=\{\bm{s}_{1},\bm{s}_{2}\}$, if action 1 is executed at step $b$ in the $(n+1)$-th episode (i.e., $\bm{s}_{1}$ is selected), we can obtain that, for $\bm{s}_{1},\bm{s}_{2}\in\mathcal{S}_{2,b},\forall b\in[B].$
$\sum_{i\in[n]}\left\langle\bm{R}_{i},~{}\bm{S}_{i}\bm{\Psi}_{n+1}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle$ (14)
Taking expectations of both sides of the inequality Eq.(14), we have
$\mathbb{E}\left[\sum_{i\in[n]}\left\langle\bm{R}_{i},~{}\bm{S}_{i}\bm{\Psi}_{n+1}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle\bigg{|}\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{n}\right]$ (15)
$\sum_{i\in[n]}\left\langle\bm{\theta}^{*},~{}\bm{S}_{i}^{\top}\bm{S}_{i}\bm{\Psi}_{n+1}^{-1}(\bm{s}_{1}-\bm{s}_{2})\right\rangle$ (16)
$\min\limits_{\bm{\theta}\in\mathbb{R}^{d}}~{}~{}\frac{1}{2}\|\bm{\theta}\|_{2}^{2}$ (17)
[…]
$\leq\alpha\left[H_{n}(\bm{s}_{I_{n,b}})-H_{n}(\bm{s}_{\bar{I}_{n,b}})\right],\ n\in[N],~{}b\in[B],$
where $\bm{\phi}_{n}(\bm{s})=\bm{\Psi}_{n+1}^{-1}\bm{s}$, $H_{n}(\bm{s})=\left[\bm{s}^{\top}\bm{\phi}_{n}(\bm{s})\right]^{\frac{1}{2}}$. Eq.(17) is the final formulation of IBCB, which proposed a simple quadratic programming (QP) problem with large scale linear constraints. Several tools have been introduced to efficiently solve QP problems [44, 45, 46]. Among these approaches, OSQP [46] is the best to achieve almost linear complexity growth in problems’ dimension for large scale low-accuracy Random QP problems, which can be applied to our IBCB.
4.2.1 Algorithm Implementation
Existing inverse bandit algorithm [15] estimates parameters using an EM-like algorithm, which iteratively maximizes the likelihood estimation. However, this algorithm suffers from low computational efficiency as it requires a large number of iterations and samples to achieve accurate estimation. Furthermore, the assumption of expert updates after each action under the fully-online setting leads to a rapid increase in computation time when sampling a large number of actions, making it unsuitable for the BCB setting. In contrast, the proposed IBCB capitalizes on the BCB assumption, which provides a unified framework for both deterministic and randomized bandit policies. By formulating the inverse bandit problem as a quadratic optimization problem with a significant number of linear constraints, IBCB offers a more efficient solution through leveraging the capabilities of the OSQP optimizer. Furthermore, since solving QP problems with full constraints costs high computation resources, and to better fulfill the updating method adopted in BCB assumption, we introduce the incremental implementation of IBCB algorithm.
We provide pseudo code in the following part to describe the details of IBCB’s incremental implementation in Algorithm 2.
Algorithm 2 Incremental implementation of IBCB algorithm
0:  Initialize $\bm{\theta}_{0}=\bm{0}_{d}$, number of episodes $N$, batch size $B$, data buffer $\{\mathcal{D}_{n}\}_{n\in[N]}$, $\lambda=\lambda_{0}$, $\alpha=\alpha_{0}$, $\bm{\Phi}_{1}=\bm{O}_{d}$.
1:  for $n=1$ to $N$  do
[…]
3:     Load last episode’s selected context vectors $\bm{S}_{n}\in\mathbb{R}^{B\times d}$ from data buffer $\mathcal{D}_{n}$.
4:     Initialize current episode’s constraint set $C_{n}=\emptyset$.
5:     for $b=1$ to $B$ do
6:        Load $b$-th step of $n$-th episode’s selected item $I_{n,b}$, other candidate items $\bar{I}_{n,b}$ and all candidate context vectors $\bm{s}$.
7:        $\bm{\Phi}_{n+1}\leftarrow\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n},$
8:        $\bm{\Psi}_{n+1}\leftarrow\lambda\bm{I}_{d}+\bm{\Phi}_{n+1}$.
[…]
10:        Construct current step’s constraint element $c$ as: $c\leftarrow\{\sum_{i\in[n]}\left\langle\bm{\theta},~{}\bm{S}_{i}^{\top}\bm{S}_{i}\left[\bm{\phi}_{n}(\bm{s}_{\bar{I}_{n,b}})-\bm{\phi}_{n}(\bm{s}_{I_{n,b}})\right]\right\rangle\leq\alpha\left[H_{n}(\bm{s}_{I_{n,b}})-H_{n}(\bm{s}_{\bar{I}_{n,b}})\right]\}$, where $\bm{\phi}_{n}(\bm{s})=\bm{\Psi}_{n+1}^{-1}\bm{s}$, $H_{n}(\bm{s})=\left[\bm{s}^{\top}\bm{\phi}_{n}(\bm{s})\right]^{\frac{1}{2}}$.
11:        Update: $C_{n}\leftarrow C_{n}\cup c$.
12:     end for
[…]
14:     Solve the following optimization problem with current episode’s constraint set $C_{n}$ through OSQP optimizer:$\min\limits_{\bm{\theta}\in\mathbb{R}^{d}}\frac{1}{2}\|\bm{\theta}-\bm{\theta}_{n-1}\|_{2}^{2}\quad\text{s.t.}\ C_{n}$
[…]
16:     Use solved parameter $\bm{\theta}$ to update: $\bm{\theta}_{n}=\bm{\theta}$.
17:  end for
18:  Return Learned expert policy parameter $\bm{\theta}_{N}$ (reward parameter $\bm{\theta}^{*}$)
4.2.2 Analysis of time complexity of IBCB
After using incremental implementation, overall time complexity of IBCB solution is reduced from $O((NBM)^{\beta}d)$ (non-incremental IBCB algorithm) to $O(N(BM)^{\beta}d)$ (incremental IBCB algorithm), where $N$ is the total number of episodes, $B$ is the total number of steps in each episode, $M$ is the size of the action space, and $d$ is the context space dimension. $\beta$ is the adaptive complexity parameter corresponding to the QP problem of different constraints for OSQP. According to the experimental results and proof of OSQP, $\beta\in\left[1,2\right]$. When the quantity of constraints is fixed, $\beta$ will adjust adaptively. When there are moderate constraints ($\leq 10^{6}$ constraints), $\beta$ will tend to 1. Analysis can be found in the following part.
We states that the training time complexity of IBCB is $O((NBM)^{\beta}d)$ (non-incremental IBCB) to $O(N(BM)^{\beta}d)$ (incremental IBCB), where $N$ is the total number of episodes, $B$ is the total number of steps in each episode, $M$ is the size of the action space, and $d$ is the context space dimension. $\beta$ is the adaptive complexity parameter corresponding to the QP problem of different constraints for OSQP. According to experimental results and the proof of OSQP [4], $\beta\in\left[1,2\right]$. When the number of constraints is fixed, $\beta$ adjusts adaptively and tends to 1 when there are moderate constraints ($\leq 10^{7}$ constraints).
Proof. For any $n\in N$, i.e., in the $n$-th episode,
1. Solving matrix $\phi_{n+1}$ in Algorithm 2 in Appendix A.1 requires matrix multiplication, with a time complexity of $O(d\*B\*d)=O(Bd^{2})$.
2. Constructing $B*(M-1)$ constraints for each step in an episode with $M$ actions requires a time complexity of $O(BMd^{3})$.
3. Using the current $B*(M-1)$ constraints to solve with OSQP requires a time complexity of $O((BM)^{\beta}d)$, where $\beta$ is the adaptive complexity parameter, $\beta\in\left[1,2\right]$, close to 1 when constraints $\leq 10^{7}$ Therefore, in the incremental update training of the IBCB algorithm, the time complexity of one episode is $O(Bd^{2}+BMd^{3}+(BM)^{\beta}d)$. Considering that $d$ is usually less than 100 in practice, the basic unit time for matrix multiplication and inversion is very low. Therefore, the time complexity of one episode is $O(BM+(BM)^{\beta}d)=O((BM)^{\beta}d)$. So, in the setting of a total of $N$ episodes, the training time complexity of the incremental IBCB is $O(N(BM)^{\beta}d)$, where $\beta$ is the adaptive complexity parameter, $\beta\in\left[1,2\right]$, close to 1 when constraints $\leq 10^{7}$.
4. If the incremental learning algorithm is not used to implement IBCB, i.e. non-incremental IBCB, OSQP needs to solve $NB*(M-1)$ constraints simultaneously, so the total time complexity of the non-incremental implementation is $O((NBM)^{\beta}d)$.
4.2.3 Extension to Fairness-Aware Expert Limitation
According to [11], a fairness-aware expert selects actions based on a nonlinear transformation of their selection probabilities, such as using Softmax, to ensure fairness by not only execute the best action. However, this can significantly reduce the average reward. To address this, we’ve enhanced the top-$k$ contextual bandit recommendation algorithm from [47] with a compromise policy: choosing the final action from the top-$k$ based on normalized probabilities, which balances reward and fairness. Our IBCB can be extended to imitate the above fairness-aware expert. Specifically, during each training episode, IBCB first trains for one round to obtain the initial reward parameters $\bm{\theta}_{\tau}$ in current episode $n$. IBCB then uses $\bm{\theta}_{\tau}$ to clean up the inequality constraints in Eq. (17) as a pre-process method, removing constraints that do not meet the criteria since expert may not select the best action, and subsequently retrains for one episode based on the remaining constraints.
Specifically, considering that IBCB has a corresponding incremental learning framework, combined with the update method of the BCB for each episode, we have made a new modification to IBCB after adding fairness. Details can be found in the following pseudo code in Algorithm 3.
Algorithm 3 IBCB’s adaptation to fairness-aware expert limitation
0:  Initialize $\bm{\theta}_{0}=\bm{0}_{d}$, number of episodes $N$, batch size $B$, data buffer $\{\mathcal{D}_{n}\}_{n\in[N]}$, $\lambda=\lambda_{0}$, $\alpha=\alpha_{0}$, $\bm{\Phi}_{1}=\bm{O}_{d}$.
1:  for $n=1$ to $N$  do
[…]
3:     Load last episode’s selected context vectors $\bm{S}_{n}\in\mathbb{R}^{B\times d}$ from data buffer $\mathcal{D}_{n}$.
4:     Initialize current episode’s constraint set $C_{n}=\emptyset$.
5:     for $b=1$ to $B$ do
6:        Load $b$-th step of $n$-th episode’s selected item $I_{n,b}$, other candidate items $\bar{I}_{n,b}$ and all candidate context vectors $\bm{s}$.
7:        $\bm{\Phi}_{n+1}\leftarrow\bm{\Phi}_{n}+\bm{S}_{n}^{\top}\bm{S}_{n},$
8:        $\bm{\Psi}_{n+1}\leftarrow\lambda\bm{I}_{d}+\bm{\Phi}_{n+1}$.
[…]
10:        Construct current step’s constraint element $c$ as: $c\leftarrow\{\sum_{i\in[n]}\left\langle\bm{\theta},~{}\bm{S}_{i}^{\top}\bm{S}_{i}\left[\bm{\phi}_{n}(\bm{s}_{\bar{I}_{n,b}})-\bm{\phi}_{n}(\bm{s}_{I_{n,b}})\right]\right\rangle\leq\alpha\left[H_{n}(\bm{s}_{I_{n,b}})-H_{n}(\bm{s}_{\bar{I}_{n,b}})\right]\}$, where $\bm{\phi}_{n}(\bm{s})=\bm{\Psi}_{n+1}^{-1}\bm{s}$, $H_{n}(\bm{s})=\left[\bm{s}^{\top}\bm{\phi}_{n}(\bm{s})\right]^{\frac{1}{2}}$.
11:        Update: $C_{n}\leftarrow C_{n}\cup c$.
12:     end for
[…]
14:     Solve the following optimization problem with current episode’s constraint set $C_{n}$ through OSQP optimizer:$\min\limits_{\bm{\theta}\in\mathbb{R}^{d}}\frac{1}{2}\|\bm{\theta}-\bm{\theta}_{n-1}\|_{2}^{2}\quad\text{s.t.}\ C_{n}$
[…]
16:     Use solved parameter $\bm{\theta}$ as this episode’s initial reward parameter $\bm{\theta}_{\tau}$: $\bm{\theta}_{\tau}=\bm{\theta}$.
[…]
19:     Initialize current expisode’s unsatisfied constraint set $C_{n-}=\emptyset$.
20:     for $b=1$ to $B$ do
21:        if $\sum_{i\in[n]}\left\langle\bm{\theta}_{\tau},~{}\bm{S}_{i}^{\top}\bm{S}_{i}\left[\bm{\phi}_{n}(\bm{s}_{\bar{I}_{n,b}})-\bm{\phi}_{n}(\bm{s}_{I_{n,b}})\right]\right\rangle>\alpha\left[H_{n}(\bm{s}_{I_{n,b}})-H_{n}(\bm{s}_{\bar{I}_{n,b}})\right]$ then
22:           Construct current step’s unsatisfied constraint element $c_{-}$ as: $c_{-}\leftarrow\{\sum_{i\in[n]}\left\langle\bm{\theta}_{\tau},~{}\bm{S}_{i}^{\top}\bm{S}_{i}\left[\bm{\phi}_{n}(\bm{s}_{\bar{I}_{n,b}})-\bm{\phi}_{n}(\bm{s}_{I_{n,b}})\right]\right\rangle>\alpha\left[H_{n}(\bm{s}_{I_{n,b}})-H_{n}(\bm{s}_{\bar{I}_{n,b}})\right]\}$
23:           Update unsatisfied constraint set: $C_{n-}\leftarrow C_{n-}\cup c_{-}$.
24:        end if
25:     end for
26:     Initialize current episode’s satisfied constraint set $C_{n}^{{}^{\prime}}$ by removing unsatisfied constraints: $C_{n}^{{}^{\prime}}\leftarrow C_{n}\backslash C_{n-}$.
[…]
29:     Solve the following optimization problem with current episode’s satisfied constraint set $C_{n}^{{}^{\prime}}$ through OSQP optimizer:$\min\limits_{\bm{\theta}^{{}^{\prime}}\in\mathbb{R}^{d}}\frac{1}{2}\|\bm{\theta}^{{}^{\prime}}-\bm{\theta}_{n-1}\|_{2}^{2}\quad\text{s.t.}\ C_{n}^{{}^{\prime}}$
[…]
31:     Use solved parameter $\bm{\theta}^{{}^{\prime}}$ to update: $\bm{\theta}_{n}=\bm{\theta}^{{}^{\prime}}$.
32:  end for
33:  Return Learned expert policy parameter $\bm{\theta}_{N}$ (reward parameter $\bm{\theta}^{*}$)
5 Experiments
5.1 Baselines
5.1.1 Original Expert Algorithms with Rewards (SBUCB & SBTS)
We used SBUCB and SBTS for original expert algorithms. These algorithms’ performances are the oracle limit above all of other IL algorithms and IBCB since they have access to the rewards. SBUCB is a batched version of LinUCB [18], which updates the policy after receiving a batch of feedback data [16]. SBTS is a batched version of Linear Thompson Sampling (LinTS) [17].
5.1.2 Behavior Cloning (BC) Algorithms
We used LinearSVC [48] model (namely BaseSVC) for BC baseline, since the we assumed reward enviornment is linear in Section 3.
5.1.3 Inverse Reinforcement Learning (IRL) Algorithms
We used Bayesian Inversive Reinforcement Learning (B-IRL) [13], and current state-of-the-art Bayesian Inversive Contextual Bandits (B-ICB) [15] to serve as IRL baselines.
5.2 Datasets
We designed synthetic and real-world dataset for experiments. For both of them, we split the whole dataset into two subsets: the Online Learning (the OL-data) phase, and the Batch Testing (the BT-data) phase. Note that expert only update its policy parameters in OL-data, and do not update parameter in BT-data phase so we could design experiments on estimating expert policy’s parameters on BT-data.
5.2.1 Synthetic Dataset
@@
We set the synthetic dataset’s candidate context $\bm{s}_{I}$ with context dimension $d=10$ as a single context (behavior) for any step of an episode. At step $b$ of $n$-th episode in OL-data phase, candidate context set $\mathcal{S}_{n,b}\subseteq{\mathbb{R}^{d}}$ was drawn from a Gaussian distribution $\mathcal{N}(\mu_{s}\bm{1}_{d},\sigma_{s}^{2}\bm{I}_{d})$, where the means of candidate contexts were $\mu_{s}\in\left[1:-0.4:-2.6\right]$, and the standard deviation was $\sigma_{s}^{2}={0.05}^{2}$. To simulate out-of-distribution(OOD) scenario, the mean of the first candidate context in the BT-data phase was set to $1.4$. We set the observed reward of a given candidate context $\bm{s}_{I_{n,b}}\in\mathcal{S}_{n,b}$ at step $b$ in $n$-th episode as a sigmoid function $\mathrm{sigmoid}(\left\langle\bm{w}_{\mathrm{reward}},\bm{s}_{I_{n,b}}\right\rangle)$, where each element of $\bm{w}_{\mathrm{reward}}\in\mathbb{R}^{d}$ was sampled from a Gaussian distribution as $\mathcal{N}(0.1,{0.01}^{2})$. To test the robustness of the model and simulate random selections of actions made by users, we also added a zero-mean Gaussian noise $\mathrm{std}\sim\mathcal{N}(0,\epsilon^{2}),\epsilon\in\{0.03,0.07,0.10\}$ in the reward, so final reward was set to be $\mathrm{sigmoid}(\left\langle\bm{w}_{\mathrm{reward}},\bm{s}_{I_{n,b}}\right\rangle)+\mathrm{std}$. We also made adjustments for the ablation study period. Contradictory data was introduced as duplicated data: we cloned the first k episodes’ candidate context sets for the second k episode’s candidate context sets, and design k with the name of duplicated quantity $\mathrm{dup}\in\left[0:1:9\right]$. Total number of conflict executed actions (e.g. expert select action 1 at step $b$ of $n$-th episode, but select action 2 at step $b$ of $(n+k)$-th episode.) of the contradictory data was ascending when we increased $\mathrm{dup}$. We also controlled the training rate for all baselines and IBCB: we set the end episode of the exposed expert’s behavioral evolution history logs with constraint end rate $\mathrm{ce\_rate}\in\left[0,1\right]$ ($e.g.$ if $\mathrm{ce\_rate}=0.5$. we used the first $\left\lfloor 0.5N\right\rfloor$ episodes for baselines and IBCB training).
In a nutshell, we generated the synthetic data for both OL-data and BT-data as follows: numbers of episodes $N=20$, batch size $B=1000$, number of candidate contexts (actions) is $M=10$ at any single step of an episode.
5.2.2 Real-World Dataset
@@
We used MovieLens 100K  [^1] https://grouplens.org/datasets/movielens/100k/ (ML-100K) dataset as our real-world experiment dataset. , which collected 5-star movie ratings from a movie recommendation service, including 943 unique users and 1682 unique movies (items). ML-100K dataset also provided various side information of both users and items. We split ML-100K dataset into OL-Data and BT-data according to the timestamps, where first half was treated as OL-data and the last half was BT-data. For reward setting, we treated the 4-5 star ratings as positive feedback(labeled with 1), and 1-3 star rating as negative feedback (labeled with 0). Through feature embedding [49] and principal component analysis (PCA), dimension of each [user, item] pair (candidate context, i.e., behavior) was reduced to 50, $i.e.,\ d=50$. For the candidate context set, we retained the original item for each interaction and then randomly sampled $9$ extra items from the entire item set (exclude the original item), so the number of candidates is $M=10$. In real recommendation, user-item interaction is sparse, meaning that not all of the items can be shown to the user and have feedback [6, 7]. To overcome this issue, we trained two Matrix Factorization (MF) [50] models using OL-data (denotes MF-OL) and BT-data (denotes MF-BT) respectively for different baselines and IBCB model to serve as the simulated real environment. AUCs of the two MF models were both over $90\%$, which assures that the simulated environment can provide nearly realistic feedbacks of users. At step $b$ of $n$-th episode of both OL-data and BT-data phase, the simulated environment (MF-OL/MF-BT) received a selected candidate context $\bm{s}_{I_{n,b}}$ and a context set $\mathcal{S}_{n,b}$ consisted of all candidate contexts in this step, getting all candidate context’s reward and sort out the candidate context $\bm{s}_{I_{\mathrm{max}}}$ with max reward $r_{I_{\mathrm{max}}}$. If selected candidate context $\bm{s}_{I_{n,b}}$ do not have the biggest reward compared to $\bm{s}_{I_{\mathrm{max}}}$, $i.e.,r_{I_{n,b}}<r_{I_{\mathrm{max}}}$, reward feedback would be directly set as 0, otherwise would be set as $r_{I_{\mathrm{max}}}$. If $r_{I_{\mathrm{max}}}>0.5$, then the final reward feedback would be 1, otherwise would be 0. Adjustments for the ablation study period was also made similar with the synthetic dataset. Since duplicated data will not appeared in a large scale in real world, we just controlled the training rate similarly as that in synthetic dataset’s setting.
@@
Finally, for OL-data in real-world dataset, we set numbers of episodes $N=20$, batch size $B=1000$, and for BT-data, we set $N=50$, $B=1000$ for precise measurement, considering real-world data’s noise. Number of candidate contexts (actions) is $M=10$ at any single step of an episode in both OL-data and BT-data phase.
5.3 Evaluation Protocol
5.3.1 Online Train Log Fitness (OL-Fitness)
OL-Fitness measures the alignment of various algorithms’ learned parameters with the true reward environment’s parameters. It involves using the learned parameters from IL algorithms (including IBCB) or backbone bandit algorithms to simulate a reward environment, where an expert interacts to gather data for online learning. The actions taken by the expert in this simulated environment are compared against those taken in the true reward environment, calculated as $\frac{1}{NB}\sum_{n=1}^{N}\sum_{b=1}^{B}\mathbb{I}(I_{n,b-\mathrm{estimated}}=I_{n,b-\mathrm{true}})$. Here, $I_{n,b-\mathrm{estimated}}$ denotes the action index from the estimated parameters in step $b$ of episode $n$ during the OL-data phase, and $I_{n,b-\mathrm{true}}$ is the action index from the true reward parameters under the same conditions. OL-data’s candidate context sets, sequences and reward parameters keep same during the comparison by using fixed seeds. Note that BC algorithms cannot obtain reward parameters, and thus were not included in this comparison scope. But in dup-test period, we used ‘train log behaviors fitness’ to evaluate whether BC algorithm suffers from contradictory data. A higher OL-Fitness indicates a closer match between the learned parameters and the true reward parameters.
5.3.2 Batch Test Log Fitness (BT-Fitness)
BT-Fitness measures the alignment of parameters/models from various algorithms with those of the final expert policy. It is calculated by comparing the actions executed by IL algorithms against those of the expert policy during the BT-data phase, using the formula $\frac{1}{NB}\sum_{n=1}^{N}\sum_{b=1}^{B}\mathbb{I}(I_{n,b-\mathrm{IL}}=I_{n,b-\mathrm{expert}})$. Here, $I_{n,b-\mathrm{IL}}$ represents the action index executed by IL algorithms (including IBCB), while $I_{n,b-\mathrm{expert}}$ is the action index from the expert policy in step $b$ of episode $n$. BT-data’s candidate context sets, sequences and reward parameters keep same during the comparison by using fixed seeds, and expert do not update its policy parameters in BT-data phase. Higher BT-Fitness indicates greater similarity to the expert policy’s parameters.
5.3.3 Batch Test Average Reward (BT-AR)
measures the reward performance of IL algorithms (including IBCB) against expert algorithms during the BT-data phase. It calculates average reward from executed actions after interacting with the BT-data environment, denoted as $\frac{1}{NB}\sum_{n=1}^{N}\sum_{b=1}^{B}R_{n,b}$. $R_{n,b}$ is binary reward for actions at step $b$ of episode $n$. Higher BT-AR indicates more effective actions in the BT-data environment, suggesting superior reward performance by the algorithm.
5.3.4 Cumulative Fairness Regret (CFR)
CFR measures the similarity in fairness between the policies of various IL algorithms (including IBCB) and the bandit policy after the final episode (i.e., the experienced expert) with fairness optimization. Following definitions of fairness regret in [11, 12] , we define CFR as $\text{CFR}=\sum_{n=1}^{N}\sum_{b=1}^{B}{(D_{\mathrm{JS}}(\pi_{\mathrm{IL}}(\mathcal{S}_{n,b})||\pi_{\mathrm{expert}}(\mathcal{S}_{n,b}))}+|R_{n,b-\mathrm{IL}}-R_{n,b-\mathrm{expert}}|),\ R_{n,b-\mathrm{IL}},R_{n,b-\mathrm{expert}}\in\{0,1\}$, where $D_{\mathrm{JS}}(P||Q)$ represents the Jensen-Shannon divergence [51] between distributions $P$ and $Q$. $\pi_{\mathrm{IL}}(\mathcal{S}_{n,b})$ refers to the policy obtained by various IL algorithms (including IBCB), and $\pi_{\mathrm{expert}}(\mathcal{S}_{n,b})$ refers to the bandit policy after the final episode (i.e., the experienced expert). Both of the above are the normalized probabilities of each context which can be selected in the candidate context list $\mathcal{S}_{n,b}$ at step $b$ of episode $n$ in BT-data phase. $R_{n,b-\mathrm{IL}}$ and $R_{n,b-\mathrm{expert}}$ are executed action’s observed reward at step $b$ of the $n$-th episode in BT-data phase from learned model and final expert respectively. It should be noted that CFR is only used in the BT-data phase, that is, when the parameters of all models have been fixed. A lower CFR indicates a closer similarity to the fairness of the final expert policy, that is, a more similar performance to the expert policy with fairness optimization.
5.4 Training Details
We run all of the experiments on the server equipped with 16 Intel Xeon E5-2630 v4@2.20GHz cores, and during training, all of the baselines and IBCB only uses the CPU. Note that all baselines and IBCB were training on expert behavioral evolution history of Online Learning. Detailed settings of baselines and IBCB are shown as follows.
5.4.1 Original Expert Policy Settings